How do you solve 5r-s=5,-4r+5s=17 for substitution property???
Answer by rfer(16322)

5r-s=5 s=5r-5
-4r+5s=17
-4r+5(5r-5)=17
21r=42
r=2
-----------------
-4r+5s=17
-4*2+5s=17
5s=25
s=5
----------
5*2-5=5